# Motion-Aware Transformer for Multi-Object Tracking

### v2

@@ 1 Introduction
Multi-object tracking (MOT) aims to detect objects and maintain their identities across frames in a video. Recent advances in query-based detection have reshaped this task. The introduction of DETR [carion2020end] marked a turning point, as it replaced handcrafted post-processing with an end-to-end formulation. This query-based design allows the network to propagate information across frames, making end-to-end training feasible and effective.
@@
Building on this paradigm, TrackFormer [meinhardt2022trackformer] extended DETR to MOT by propagating detection queries forward as track queries, and concatenating them with newly initialized detection queries for subsequent frames. This formulation enabled consistent learning of track queries without manual association rules. Later, MOTR [zeng2022motr] introduced a Temporal Aggregation Network (TAN) to update track queries between frames, alleviating issues of overfitting and improving performance on challenging datasets such as DanceTrack. These advances have established query propagation as a dominant strategy in end-to-end MOT.
@@
In addition to introducing MATR, we strengthen the MOTR baseline with improved training strategies and updated DETR-family components such as DAB-DETR [liu2022dabdetr]. This provides a stronger foundation for comparison and ensures that the observed improvements stem from the proposed motion-aware design. Comprehensive experiments across three benchmarks validate our approach. On the DanceTrack dataset, MATR improves HOTA by more than 9 points over MOTR without additional data, and reaches a state-of-the-art HOTA of 71.3 when supplementary data are included. On SportsMOT, MATR achieves a HOTA of 72.2, establishing a new state of the art without relying on extra datasets. Finally, in multi-category tracking on BDD100k, MATR obtains 54.7 mTETA and 41.6 mHOTA, surpassing previous methods under the same training conditions.
@@ 2 Related Work
Tracking-by-Detection remains the most common paradigm for multi-object tracking. These methods typically adopt a two-step pipeline: first detecting bounding boxes in each frame, then associating detections across frames to construct trajectories. Performance is therefore heavily dependent on the accuracy of the detection stage. SORT [bewley2016simple] exemplifies this approach by using Hungarian algorithm [kuhn1955hungarian] for association and a Kalman filter [welch1995kalman] to predict object positions. Detections are matched to predicted boxes using intersection-over-union (IoU) scores. DeepSORT [Wojke2017simple] extends SORT by incorporating appearance features extracted by a learned embedding network, which are compared via cosine distance to improve association robustness. JDE [wang2020towards], FairMOT [zhang2021fairmot], and Unicorn [yan2022towards] jointly train detection and appearance embeddings in a unified framework. ByteTrack [zhang2022bytetrack], based on a YOLOX detector, achieves state-of-the-art performance by considering both high- and low-confidence detections. BoT-SORT [aharon2022botsort] enhances SORT by refining the Kalman filter state, introducing camera-motion compensation, and integrating Re-ID features. Recent efforts incorporate transformers into the association. TransMOT [chu2023transmot] and GTR [zhou2022global] use spatial-temporal transformers to model interactions between instances and aggregate historical context. OC-SORT [cao2023observation] improves upon SORT by re-identifying and rehabilitating lost tracks.
Query Propagation has emerged as a popular alternative with the rise of DETR-family models, end-to-end approach for multi-object tracking. These methods maintain query embeddings across frames to track individual objects, eliminating the need for handcrafted post-processing. Tracktor++ [bergmann2019tracktor] pioneered this idea by leveraging an R-CNN [girshick2014rcnn] regression head to re-localize bounding boxes across frames. TrackFormer [meinhardt2022trackformer] extended Deformable DETR [zhu2021deformable] to propagate detection queries as track queries, while MOTR [zeng2022motr] introduced a Temporal Aggregation Network (TAN) to continuously update track queries. MeMOT [cai2022memot] improved query propagation by employing both short- and long-term memory banks, while TransTrack [sun2020transtrack] propagated queries across frames for continuous detection. Other extensions include P3AFormer [zhao2022tracking], which integrates optical-flow-guided feature propagation, and CO-MOT [yan2023bridging], which introduces one-to-set label assignment and a COopetition Label Assignment (COLA) strategy to enhance detection performance. MeMOTR [gao2023memotr] further improved association by incorporating historical frame information. MOTRv2 [zhang2023motrv2] integrated a YOLOX detector to strengthen detection with external proposals. MOTRv3 [yu2023motrv3] introduced RFS to mitigate label insufficiency, Pseudo Label from a pretrained 2D detector, and Query Grouping with Disturbance.
@@ 3.1 Baseline Improvement
Since the introduction of MOTR, which was built on Deformable DETR [zhu2021deformable], newer DETR-family frameworks such as DAB-Deformable DETR [liu2022dabdetr] have demonstrated improved detection performance. To establish a stronger baseline, we adopt the bounding box propagation mechanism, where bounding boxes are propagated as positional encodings for queries. Specifically, bounding boxes are initialized randomly as $N_{\det}\times[x_{0},y_{0},h_{0},w_{0}]$, where $N_{\det}$ is the number of detection queries, and iteratively refined by successive decoder layers as $N_{\det}\times[\Delta x,\Delta y,\Delta h,\Delta w]$. It is important to note that we use only the box propagation strategy from [zhu2021deformable], rather than adopting the full DAB model. Fully incorporating DAB increases parameters from 42M to 49M, but unexpectedly reduces HOTA from 69.4 to 67.6. We observed that larger models generated more bounding boxes, yet overall performance degraded, likely due to overfitting in the absence of additional augmentation.
@@ 3.3 Motion-aware Transformer
Unlike prior works such as MOTR [zeng2022motr, zhang2023motrv2] which update track query features solely through self-attention, we argue that this design suffers from three key limitations: 1) without explicit supervision and guidance, track queries do not learn motion effectively; 2) self-attention alone fails to properly integrate the previous track queries $Q_{\text{trk}}^{i-1}$ with features from the current frame, which negatively impacts the subsequent Transformer Decoder; and 3) leaving positional embeddings unchanged can cause a mismatch between feature and positional information, further impairing training.
@@ 3.4 Inference Details
The inference procedure largely follows MOTR [zeng2022motr], with adaptations to accommodate MAT. Since ground truth is unavailable during inference, detection queries are not filtered before being processed by MAT. To handle occlusions, we filter the outputs of the Transformer Decoder. If the confidence score of a tracked object falls below the threshold $\tau_{\text{trk}}$ in the current frame, we retain its embedding temporarily as an inactive trajectory. If its confidence remains below the threshold for more than $\mathcal{T}_{\text{miss}}$ consecutive frames, the inactive trajectory is permanently removed. Similarly, detection results are discarded if their confidence falls below $\tau_{\det}$. In all experiments, we set the thresholds to $\tau_{\det}=0.7$, $\tau_{\text{trk}}=0.5$, and $\mathcal{T}_{\text{miss}}=25$, unless otherwise specified.
@@ 4.1 Dataset and Metrics
We evaluate our approach on three challenging benchmarks: DanceTrack [sun2022dancetrack], SportsMOT [cui2023sportsmot], and BDD100k [bdd100k].
@@
FairMOT [zhang2021fairmot] / 39.7 66.7 23.8 82.2 40.8
CenterTrack [zhou2020tracking] / 41.8 78.1 22.6 86.8 35.7
TraDeS [wu2021track] / 43.3 74.5 25.4 86.2 41.2
ByteTrack [zhang2022bytetrack] / 47.7 71.0 32.1 89.6 53.9
GTR [zhou2022global] / 48.0 72.5 31.9 84.7 50.3
QDTrack [pang2021quasi] / 54.2 80.1 36.8 87.7 50.4
OC-SORT [cao2023observation] / 55.1 80.3 38.3 92.0 54.6
C-BIoU [yang2023hard] / 60.6 81.3 45.4 91.6 61.6
TransTrack [sun2020transtrack] / 45.5 75.9 27.5 88.4 45.2
MOTR [zeng2022motr] 40 54.2 73.5 40.2 79.7 51.5
MOTRv3-R50 [yu2023motrv3] 40 63.9 76.7 53.5 86.8 67.2
MeMOTR [gao2023memotr] 48 68.5 80.5 58.4 89.9 71.2
CO-MOT [yan2023bridging] 40 69.4 82.1 58.9 91.2 71.9
@@
MOTRv2 [zhang2023motrv2] 94 69.9 83.0 59.0 91.9 71.7
MOTRv3 [yu2023motrv3] 103 70.4 83.8 59.3 92.9 72.3
@@
For evaluation, we primarily use the Higher Order Tracking Accuracy (HOTA) metric [luiten10higher], which provides a balanced assessment of tracking by decomposing results into detection accuracy (DetA) and association accuracy (AssA). We additionally report the widely used MOTA and IDF1 metrics to give a broader view of detection and identification performance. For multi-category scenarios, we adopt Track Every Thing Accuracy (TETA) [li2022tracking], which includes sub-metrics for localization accuracy (LocA) and association accuracy (AssocA). TETA has been shown to provide a more reliable and comprehensive evaluation than traditional metrics, especially on large-scale multi-class datasets such as BDD100k.
@@ 4.2 Implementation Details
We build MATR upon MOTR [zeng2022motr] and employ several data augmentation strategies, including random resizing and random cropping. The shorter side of each input image is resized to 800 pixels, with the longer side restricted to a maximum of 1536 pixels. Following [zeng2022motr], we use ResNet50 as the backbone network. Additionally, we evaluate a larger backbone, Swin-Tiny [liu2021swin], to provide a more comprehensive comparison. For hyperparameters, the detection queries $N_{\det}$ is set to 300, the sequence length $S$ is fixed at 5, and the dropout probability for track queries is 0.1. The batch size is 1 per GPU, with each batch containing a video clip of $S$ frames sampled at random intervals from 1 to 10 frames. We adopt the AdamW optimizer, using an initial learning rate of $2\times 10^{-4}$ for ResNet50 and $1\times 10^{-4}$ for Swin. Tracked targets with IoU scores below the threshold $\tau_{\text{IoU}}=0.5$ are filtered out. These training settings remain consistent across all experiments. For the CrowdHuman dataset, we follow CenterTrack [zhou2020tracking] and apply random spatial shifts to generate pseudo trajectories when including it in training. The overall MATR loss is defined as:
@@
Following standard practice in MOTR, CO-MOT, and MeMOTR [zhang2023motrv2, yu2023motrv3, yan2023bridging], we combine DanceTrack training data with CrowdHuman. Training lasts 20 epochs with a learning rate decay of 0.1 applied every 8 epochs, requiring about 2.5 days in total. For SportsMOT, training follows the same setup as DanceTrack but excludes additional datasets. For BDD100k, the reduced input size and classification head modification match the settings used in prior end-to-end baselines, ensuring comparability. We also follow the SportsMOT setup, except training runs for 13 epochs with a learning rate decay at epoch 11, takes approximately 7 days.
@@ 4.3 Ablation Study
Since MAT predicts future positions in advance, it shares some conceptual similarity with the Kalman Filter [welch1995kalman] (KLF), a classic prediction mechanism used in traditional tracking-by-detection methods. We therefore include KLF as a baseline in Table 3. The results indicate that while association accuracy (AssA) increases whenever a prediction mechanism is applied, detection accuracy (DetA) drops sharply when using KLF, indicating that its poor prediction quality prevents high-quality detections from being matched. This shows that KLF’s prediction accuracy is insufficient for end-to-end MOT. The same limitation has long been observed in tracking-by-detection pipelines: their methods typically achieve higher detection accuracy, but the low-quality predictions from KLF often prevent high-quality detections from being matched during association, ultimately degrading performance. In end-to-end methods, this issue is further magnified - poor predictions from KLF directly propagate into the Transformer Decoder, causing a substantial decline in detection accuracy.
@@ 4.4 Comparison on DanceTrack
MATR delivers significant improvements in both association accuracy and overall tracking performance. Table 1 reports a comparison between MATR and other SOTA methods on the DanceTrack test set. Our method achieves a new SOTA HOTA score of 71.3, with particularly strong association metrics, including 61.6 AssA and 75.3 IDF1. Due to the inherent limitations of linear motion prediction, tracking-by-detection methods often excel at detection accuracy but fall short in association. For example, OC-SORT [cao2023observation] achieves excellent detection accuracy (DetA = 80.3) but struggles with association in complex motion scenarios (AssA = 38.3). MATR improves both detection and association metrics simultaneously, highlighting the benefit of explicitly modeling motion to reduce query collisions.
For fairness, we also report MATR results with a ResNet50 backbone aligned with [zeng2022motr, gao2023memotr, yan2023bridging] , denoted as MATR-R50. We additionally include MOTRv3-R50 [yu2023motrv3] for reference. Compared with other ResNet50-based methods, MATR achieves a HOTA score of 69.4, matching CO-MOT. Notably: (i) MATR improves HOTA by 15.2 points over MOTR, together with substantial gains across all other metrics; (ii) MATR surpasses MeMOTR by 0.9 HOTA and consistently achieves stronger association performance, without relying on merging strategies across historical frames; (iii) MATR matches CO-MOT’s HOTA performance despite not employing explicit detection enhancements, trailing slightly in DetA (-0.6) but outperforming in AssA (+0.2) and IDF1 (+0.7). These findings demonstrate that the consistent training strategies introduced by MAT not only boost detection accuracy but also yield even greater improvements in association quality.
@@
We further compare MATR with larger and more complex models such as MOTRv2/3 [zhang2023motrv2, yu2023motrv3]. MATR achieves a HOTA of 71.3, surpassing MOTRv3 by 0.9 points, while maintaining an elegant and concise end-to-end structure. In contrast, MOTRv2 incorporates YOLOX as an external detector, and MOTRv3 employs multiple complex strategies. Despite their complexity, these methods rely on significantly larger backbones (94M and 103M parameters, respectively), whereas MATR achieves superior performance with a smaller Swin-Tiny backbone (43M parameters). Although MATR trails MOTRv3 slightly in DetA (-1.2) and MOTA (-1.0) due to the absence of specialized detection enhancement techniques, it achieves state-of-the-art association results (+2.3 AssA and +3.0 IDF1). This confirms the robustness and effectiveness of explicitly addressing query collisions.
@@
Figure 6: Qualitative comparison between MOTR and MATR. Case A shows that MATR successfully tracks a background object despite occlusion, while MOTR fails due to confusion from a foreground object. Cases B and C demonstrate that MATR maintains accurate tracking despite identity switches caused by crossing trajectories, which MOTR cannot resolve. Case D illustrates that MATR robustly tracks multiple objects in a complex scenario, while MOTR suffers from identity switches and lost trajectories after 30 frames. Note that while MATR reduces most query collisions, it can still fail under extreme occlusions or abrupt camera motion.

### v3

@@ 1 Introduction
Multi-object tracking (MOT) aims to detect objects and maintain their identities across frames in a video. Recent advances in query-based detection have reshaped this task. The introduction of DETR [6] marked a turning point, as it replaced handcrafted post-processing with an end-to-end formulation. This query-based design allows the network to propagate information across frames, making end-to-end training feasible and effective.
@@
Building on this paradigm, TrackFormer [16] extended DETR to MOT by propagating detection queries forward as track queries, and concatenating them with newly initialized detection queries for subsequent frames. This formulation enabled consistent learning of track queries without manual association rules. Later, MOTR [30] introduced a Temporal Aggregation Network (TAN) to update track queries between frames, alleviating issues of overfitting and improving performance on challenging datasets such as DanceTrack. These advances have established query propagation as a dominant strategy in end-to-end MOT.
@@
In addition to introducing MATR, we strengthen the MOTR baseline with improved training strategies and updated DETR-family components such as DAB-DETR [13]. This provides a stronger foundation for comparison and ensures that the observed improvements stem from the proposed motion-aware design. Comprehensive experiments across three benchmarks validate our approach. On the DanceTrack dataset, MATR improves HOTA by more than 9 points over MOTR without additional data, and reaches a state-of-the-art HOTA of 71.3 when supplementary data are included. On SportsMOT, MATR achieves a HOTA of 72.2, establishing a new state of the art without relying on extra datasets. Finally, in multi-category tracking on BDD100k, MATR obtains 54.7 mTETA and 41.6 mHOTA, surpassing previous methods under the same training conditions.
@@ 2 Related Work
Tracking-by-Detection remains the most common paradigm for multi-object tracking. These methods typically adopt a two-step pipeline: first detecting bounding boxes in each frame, then associating detections across frames to construct trajectories. Performance is therefore heavily dependent on the accuracy of the detection stage. SORT [3] exemplifies this approach by using Hungarian algorithm [11] for association and a Kalman filter [22] to predict object positions. Detections are matched to predicted boxes using intersection-over-union (IoU) scores. DeepSORT [23] extends SORT by incorporating appearance features extracted by a learned embedding network, which are compared via cosine distance to improve association robustness. JDE [21], FairMOT [32], and Unicorn [25] jointly train detection and appearance embeddings in a unified framework. ByteTrack [31], based on a YOLOX detector, achieves state-of-the-art performance by considering both high- and low-confidence detections. BoT-SORT [1] enhances SORT by refining the Kalman filter state, introducing camera-motion compensation, and integrating Re-ID features. Recent efforts incorporate transformers into the association. TransMOT [7] and GTR [36] use spatial-temporal transformers to model interactions between instances and aggregate historical context. OC-SORT [5] improves upon SORT by re-identifying and rehabilitating lost tracks.
Query Propagation has emerged as a popular alternative with the rise of DETR-family models, end-to-end approach for multi-object tracking. These methods maintain query embeddings across frames to track individual objects, eliminating the need for handcrafted post-processing. Tracktor++ [2] pioneered this idea by leveraging an R-CNN [10] regression head to re-localize bounding boxes across frames. TrackFormer [16] extended Deformable DETR [37] to propagate detection queries as track queries, while MOTR [30] introduced a Temporal Aggregation Network (TAN) to continuously update track queries. MeMOT [4] improved query propagation by employing both short- and long-term memory banks, while TransTrack [20] propagated queries across frames for continuous detection. Other extensions include P3AFormer [34], which integrates optical-flow-guided feature propagation, and CO-MOT [26], which introduces one-to-set label assignment and a COopetition Label Assignment (COLA) strategy to enhance detection performance. MeMOTR [9] further improved association by incorporating historical frame information. MOTRv2 [33] integrated a YOLOX detector to strengthen detection with external proposals. MOTRv3 [28] introduced RFS to mitigate label insufficiency, Pseudo Label from a pretrained 2D detector, and Query Grouping with Disturbance.
@@ 3.1 Baseline Improvement
Since the introduction of MOTR, which was built on Deformable DETR [37], newer DETR-family frameworks such as DAB-Deformable DETR [13] have demonstrated improved detection performance. To establish a stronger baseline, we adopt the bounding box propagation mechanism, where bounding boxes are propagated as positional encodings for queries. Specifically, bounding boxes are initialized randomly as $N_{\det}\times[x_{0},y_{0},h_{0},w_{0}]$, where $N_{\det}$ is the number of detection queries, and iteratively refined by successive decoder layers as $N_{\det}\times[\Delta x,\Delta y,\Delta h,\Delta w]$. It is important to note that we use only the box propagation strategy from [37], rather than adopting the full DAB model. Fully incorporating DAB increases parameters from 42M to 49M, but unexpectedly reduces HOTA from 69.4 to 67.6. We observed that larger models generated more bounding boxes, yet overall performance degraded, likely due to overfitting in the absence of additional augmentation.
@@ 3.3 Motion-aware Transformer
Unlike prior works such as MOTR [30, 33] which update track query features solely through self-attention, we argue that this design suffers from three key limitations: 1) without explicit supervision and guidance, track queries do not learn motion effectively; 2) self-attention alone fails to properly integrate the previous track queries $Q_{\text{trk}}^{i-1}$ with features from the current frame, which negatively impacts the subsequent Transformer Decoder; and 3) leaving positional embeddings unchanged can cause a mismatch between feature and positional information, further impairing training.
@@ 3.4 Inference Details
The inference procedure largely follows MOTR [30], with adaptations to accommodate MAT. Since ground truth is unavailable during inference, detection queries are not filtered before being processed by MAT. To handle occlusions, we filter the outputs of the Transformer Decoder. If the confidence score of a tracked object falls below the threshold $\tau_{\text{trk}}$ in the current frame, we retain its embedding temporarily as an inactive trajectory. If its confidence remains below the threshold for more than $\mathcal{T}_{\text{miss}}$ consecutive frames, the inactive trajectory is permanently removed. Similarly, detection results are discarded if their confidence falls below $\tau_{\det}$. In all experiments, we set the thresholds to $\tau_{\det}=0.7$, $\tau_{\text{trk}}=0.5$, and $\mathcal{T}_{\text{miss}}=25$, unless otherwise specified.
@@ 4.1 Dataset and Metrics
We evaluate our approach on three challenging benchmarks: DanceTrack [19], SportsMOT [8], and BDD100k [29].
@@
FairMOT [32] / 39.7 66.7 23.8 82.2 40.8
CenterTrack [35] / 41.8 78.1 22.6 86.8 35.7
TraDeS [24] / 43.3 74.5 25.4 86.2 41.2
ByteTrack [31] / 47.7 71.0 32.1 89.6 53.9
GTR [36] / 48.0 72.5 31.9 84.7 50.3
QDTrack [17] / 54.2 80.1 36.8 87.7 50.4
OC-SORT [5] / 55.1 80.3 38.3 92.0 54.6
C-BIoU [27] / 60.6 81.3 45.4 91.6 61.6
TransTrack [20] / 45.5 75.9 27.5 88.4 45.2
MOTR [30] 40 54.2 73.5 40.2 79.7 51.5
MOTRv3-R50 [28] 40 63.9 76.7 53.5 86.8 67.2
MeMOTR [9] 48 68.5 80.5 58.4 89.9 71.2
CO-MOT [26] 40 69.4 82.1 58.9 91.2 71.9
@@
MOTRv2 [33] 94 69.9 83.0 59.0 91.9 71.7
MOTRv3 [28] 103 70.4 83.8 59.3 92.9 72.3
@@
For evaluation, we primarily use the Higher Order Tracking Accuracy (HOTA) metric [15], which provides a balanced assessment of tracking by decomposing results into detection accuracy (DetA) and association accuracy (AssA). We additionally report the widely used MOTA and IDF1 metrics to give a broader view of detection and identification performance. For multi-category scenarios, we adopt Track Every Thing Accuracy (TETA) [12], which includes sub-metrics for localization accuracy (LocA) and association accuracy (AssocA). TETA has been shown to provide a more reliable and comprehensive evaluation than traditional metrics, especially on large-scale multi-class datasets such as BDD100k.
@@ 4.2 Implementation Details
We build MATR upon MOTR [30] and employ several data augmentation strategies, including random resizing and random cropping. The shorter side of each input image is resized to 800 pixels, with the longer side restricted to a maximum of 1536 pixels. Following [30], we use ResNet50 as the backbone network. Additionally, we evaluate a larger backbone, Swin-Tiny [14], to provide a more comprehensive comparison. For hyperparameters, the detection queries $N_{\det}$ is set to 300, the sequence length $S$ is fixed at 5, and the dropout probability for track queries is 0.1. The batch size is 1 per GPU, with each batch containing a video clip of $S$ frames sampled at random intervals from 1 to 10 frames. We adopt the AdamW optimizer, using an initial learning rate of $2\times 10^{-4}$ for ResNet50 and $1\times 10^{-4}$ for Swin. Tracked targets with IoU scores below the threshold $\tau_{\text{IoU}}=0.5$ are filtered out. These training settings remain consistent across all experiments. For the CrowdHuman dataset, we follow CenterTrack [35] and apply random spatial shifts to generate pseudo trajectories when including it in training. The overall MATR loss is defined as:
@@
Following standard practice in MOTR, CO-MOT, and MeMOTR [33, 28, 26], we combine DanceTrack training data with CrowdHuman. Training lasts 20 epochs with a learning rate decay of 0.1 applied every 8 epochs, requiring about 2.5 days in total. For SportsMOT, training follows the same setup as DanceTrack but excludes additional datasets. For BDD100k, the reduced input size and classification head modification match the settings used in prior end-to-end baselines, ensuring comparability. We also follow the SportsMOT setup, except training runs for 13 epochs with a learning rate decay at epoch 11, takes approximately 7 days.
@@ 4.3 Ablation Study
Since MAT predicts future positions in advance, it shares some conceptual similarity with the Kalman Filter [22] (KLF), a classic prediction mechanism used in traditional tracking-by-detection methods. We therefore include KLF as a baseline in Table 3. The results indicate that while association accuracy (AssA) increases whenever a prediction mechanism is applied, detection accuracy (DetA) drops sharply when using KLF, indicating that its poor prediction quality prevents high-quality detections from being matched. This shows that KLF’s prediction accuracy is insufficient for end-to-end MOT. The same limitation has long been observed in tracking-by-detection pipelines: their methods typically achieve higher detection accuracy, but the low-quality predictions from KLF often prevent high-quality detections from being matched during association, ultimately degrading performance. In end-to-end methods, this issue is further magnified - poor predictions from KLF directly propagate into the Transformer Decoder, causing a substantial decline in detection accuracy.
@@ 4.4 Comparison on DanceTrack
MATR delivers significant improvements in both association accuracy and overall tracking performance. Table 1 reports a comparison between MATR and other SOTA methods on the DanceTrack test set. Our method achieves a new SOTA HOTA score of 71.3, with particularly strong association metrics, including 61.6 AssA and 75.3 IDF1. Due to the inherent limitations of linear motion prediction, tracking-by-detection methods often excel at detection accuracy but fall short in association. For example, OC-SORT [5] achieves excellent detection accuracy (DetA = 80.3) but struggles with association in complex motion scenarios (AssA = 38.3). MATR improves both detection and association metrics simultaneously, highlighting the benefit of explicitly modeling motion to reduce query collisions.
For fairness, we also report MATR results with a ResNet50 backbone aligned with [30, 9, 26] , denoted as MATR-R50. We additionally include MOTRv3-R50 [28] for reference. Compared with other ResNet50-based methods, MATR achieves a HOTA score of 69.4, matching CO-MOT. Notably: (i) MATR improves HOTA by 15.2 points over MOTR, together with substantial gains across all other metrics; (ii) MATR surpasses MeMOTR by 0.9 HOTA and consistently achieves stronger association performance, without relying on merging strategies across historical frames; (iii) MATR matches CO-MOT’s HOTA performance despite not employing explicit detection enhancements, trailing slightly in DetA (-0.6) but outperforming in AssA (+0.2) and IDF1 (+0.7). These findings demonstrate that the consistent training strategies introduced by MAT not only boost detection accuracy but also yield even greater improvements in association quality.
@@
We further compare MATR with larger and more complex models such as MOTRv2/3 [33, 28]. MATR achieves a HOTA of 71.3, surpassing MOTRv3 by 0.9 points, while maintaining an elegant and concise end-to-end structure. In contrast, MOTRv2 incorporates YOLOX as an external detector, and MOTRv3 employs multiple complex strategies. Despite their complexity, these methods rely on significantly larger backbones (94M and 103M parameters, respectively), whereas MATR achieves superior performance with a smaller Swin-Tiny backbone (43M parameters). Although MATR trails MOTRv3 slightly in DetA (-1.2) and MOTA (-1.0) due to the absence of specialized detection enhancement techniques, it achieves state-of-the-art association results (+2.3 AssA and +3.0 IDF1). This confirms the robustness and effectiveness of explicitly addressing query collisions.
@@
Figure 6: Qualitative comparison between MOTR and MATR. Case A shows that MATR successfully tracks a background object despite occlusion, while MOTR fails due to confusion from a foreground object. Cases B and C demonstrate that MATR maintains accurate tracking despite identity switches caused by crossing trajectories, which MOTR cannot resolve. Case D illustrates that MATR robustly tracks multiple objects in a complex scenario, while MOTR suffers from identity switches and lost trajectories after 30 frames.